Session: Feb 2005 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Urumqi (yang wen jun at Thu Mar 10 09:04:40 2005) Success :
The experiment was ended at UT 08:41 because of the following source is outside dish limits.

Urumqi (yang wen jun at Thu Mar 10 09:11:30 2005) Success :
Because we have not enough hard-disk, so, the data of ek020a,ez013 and eh010 were all recorded in the same hard-disk,the number of disk is cmva-010 aoed.

Torun (K. Borkowski at Thu Mar 10 10:14:41 2005) Success :
No known problems.

Onsala (Michael Lindqvist at Thu Mar 10 10:25:54 2005) Success :
No known problems.

Medicina (Zacchiroli at Thu Mar 10 10:57:40 2005) Minor failures :
Stopped from 07:29:00 till 09:15:20 for antenna problems

Hartebeesthoek (Jonathan Quick at Thu Mar 10 12:18:34 2005) Minor failures : SLEW RFI
Intermittent RFI throughout observation. Antenna drive interface failed at 08:00 UT, restarted 08:36 on calibrator but failed again at 08:50 with Declination axis brake failure. Source had set below local horizon by then so we ended experiment early.

Jodrell_Bank (Alastair Gunn at Thu Mar 10 22:28:16 2005) Minor failures :
The Lovell Telescope was used. The schedule was altered to limit the number of phase calibrator source changes; some scans will not fringe. Schedule refused to append data to disk pack at experiment start. Changed disk pack and began recording manually at 23:01:22 UT. No other known problems.

Cambridge (Alastair Gunn at Thu Mar 10 22:28:44 2005) Success :
No known problems.

Shanghai (Tao An at Fri Mar 11 02:18:00 2005) Minor failures : FS
Due to FS and MK5 problem, we lost scans from No0010 to No0013.

Effelsberg (A. Kraus at Fri Mar 11 12:56:57 2005) Success :

Noto (P. Cassaro at Mon Mar 21 10:22:31 2005) Minor failures : POIN
From 05:00 to 06:00 UT pointing problems due to high elevation.